OpenStudy (anonymous):

If one sprinter runs the 400m in 58 seconds and another can run the same distance in 60 seconds, by how much distance will the faster person win by if they race? (physics question)

OpenStudy (jack1):

so average speed of the runner who does 400 m per 58 seconds = ...? \[ \frac{400m }{ 58s } = \frac{ xm }{1s }\] solve for x

OpenStudy (jack1):

then do the same for the guy who does it on 60 sec then work out the distance the 2nd guy has travelled in 58 sec minus that figure from 400m = final answer

OpenStudy (jack1):

is that all makin sense @stephjlamb ?

OpenStudy (anonymous):

@Jack1 Kinda, my teacher gave us the equation average velocity= change in postition/change in time, would that equation work for the problem?

OpenStudy (jack1):

yeah, it's kinda already given in that the average velocity is: Runner 1 - 400m/58 seconds Runner 2 - 400m/60 seconds these are both average velocities if you want to make it easier to use, you can convert the velocities into meters travelled per 1 second, or convert to kilometers per hour or feet / day, but either way it will still be the same velocity

OpenStudy (jack1):

so back to the 2nd runner, his average velocity is 400m per 60 seconds so how far will he travel in 1 second?

OpenStudy (anonymous):

6.7 m?

OpenStudy (jack1):

\[\frac{ 400m }{ 60s } \rightarrow \frac{ x }{ 1s }\]

OpenStudy (jack1):

so 60 seconds divided by 60 = 1 second therefore 400m divided by 60 = x meters

OpenStudy (jack1):

sorry, my net is laggy, yeah, 6.7 m

OpenStudy (jack1):

so if runner 1 finished the race at 58 seconds, he's a position 400m what position is runner 2 at at 58 seconds? how far behind runner 1 is this?

OpenStudy (anonymous):

Thats ok you've been a great help! so would i just double 6.7- to 13.4 then subtract 13.4 from 400, so that would be 386.6m, and is that my answer?

OpenStudy (jack1):

perfect, spot on hey

OpenStudy (anonymous):

Thank you so much!

OpenStudy (jack1):

all good, yw

OpenStudy (anonymous):

@Jack1 I have another question, this other problem says- A wombat runs south in a straight line with an average velocity of 5 m/s for 4 minutes and then with an average velocity of 4 m/s for 3 minutes in the same direction a.) what is its total displacement- which I got as -1920 m or -1.92 km b.) What is its average velocity during this time? which the answer the worksheet gave me is -4.57 m/s and I was wondering if you knew how to calculate this?

OpenStudy (anonymous):

@Jack1

OpenStudy (jack1):

sorry was doing another q, back now

OpenStudy (anonymous):

Its ok I solved it!

OpenStudy (jack1):

sweet ...a wombat, where do u get ur maths books from, down unda store ;D

OpenStudy (anonymous):

Lol @Jack1

OpenStudy (jack1):

he he he